1. Joint Cost Allocation—Physical Units Method Board-It, Inc., produces the following types of 2 × 4...

1.

Joint Cost Allocation—Physical Units Method

Board-It, Inc., produces the following types of 2 × 4 × 10 wood boards: washed, stained, and pressure treated. These products are produced jointly until they are cut. One batch produces 45 washed boards, 35 stained boards, and 20 pressure treated boards. The joint production process costs a total of $710 per batch.

Using the physical units method, allocate the joint production cost to each product. Round your answers to two decimal places.

Joint Product Allocation
Washed $
Stained
Pressure treated
Totals $

--------------------------------------------------------------

2.

Joint Cost Allocation—Market Value at Split-off Method

Toil & Oil processes crude oil to jointly produce gasoline, diesel, and kerosene. One batch produces 3,415 gallons of gasoline, 2,732 gallons of diesel, and 1,366 gallons of kerosene at a joint cost of $12,000. After the split-off point, all products are processed further, but the estimated market price for each product at the split-off point is as follows:

Gasoline $2 per gallon
Diesel 1 per gallon
Kerosene 3 per gallon

Using the market value at split-off method, allocate the $12,000 joint cost of production to each product.

Joint Product Allocation
Gasoline $
Diesel
Kerosene
Totals $

------------------------------------------------------------------------------------

3.

Joint Cost Allocation—Weighted Average Method

Carving Creations jointly produces wood chips and sawdust used in agriculture. The wood chips and sawdust are actually by-products of the company’s core operations, but Carving Creations accounts for them just like normally produced goods because of their large volumes. One jointly produced batch yields 3,000 cubic yards of wood chips and 10,000 cubic yards of sawdust, and the estimated cost per batch is $21,400. However, the joint production of each good is not equally weighted. Management at Carving Creations estimates that for the time it takes to produce 10 cubic yards of wood chips in the joint production process, only 2 cubic yards of sawdust are produced.

Given this information, allocate the joint costs of production to each product using the weighted average method.

Joint Product Allocation
Sawdust $
Wood chips
Totals $

------------------------------

4.

Support department cost allocation—direct method

Becker Tabletops has two support departments (Janitorial and Cafeteria) and two production departments (Cutting and Assembly). Relevant details for these departments are as follows:

Support Department Cost Driver
Janitorial Department Square footage to be serviced
Cafeteria Department Number of employees
Janitorial
Department
Cafeteria
Department
Cutting
Department
Assembly
Department
Department costs $280,000    $170,000 $1,510,000    $670,000
Square feet 46 5,300 2,100 4,900
Number of employees 12 3    18 2   

Allocate the support department costs to the production departments using the direct method.

Cutting
Department
Assembly
Department
Janitorial Department cost allocation $ $
Cafeteria Department cost allocation $ $
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Board-It, Inc. A B C=A/ Sum of A D=B*C
Final Answer Qty. % Joint Cost Allocated cost
Washed                  45.00 45.00%                   710.00               319.50
Stained                  35.00 35.00%                   710.00               248.50
Pressure Treated                  20.00 20.00%                   710.00               142.00
Total               100.00               710.00
Toil & Oil A B C=A*B D=C/ Sum of C E F=E*D
Final Answer Qty. Sell Price Market Value % Joint Cost Allocated cost
Gasoline            3,415.00                2.00               6,830.00 50.00%         12,000.00      6,000.00
Diesel            2,732.00                1.00               2,732.00 20.00%         12,000.00      2,400.00
Kerosene            1,366.00                3.00               4,098.00 30.00%         12,000.00      3,600.00
Total             13,660.00 12,000.00
Carving Creations A B C=A/B D=C/ Sum of C E F=E*D
Final Answer Qty. Weight Time Taken % Joint Cost Allocated cost
Sawdust          10,000.00                2.00               5,000.00 94.34%         21,400.00 20,188.68
Woodchips            3,000.00              10.00                   300.00 5.66%         21,400.00      1,211.32
Total               5,300.00 21,400.00
Becker Tabletops
Direct Method Janitorial Cafeteria Cutting Assembly Total
Square Feet               2,100.00            4,900.00            7,000.00
Square Feet % 30.00% 70.00% 100.00%
Number of Employees                     18.00                   2.00                 20.00
Number of Employees % 90.00% 10.00% 100.00%
Total costs        280,000.00    170,000.00        1,510,000.00         67,000.00 2,027,000.00
Janitorial cost to be allocated- Square Feet %      (280,000.00)                     -               84,000.00       196,000.00                        -  
Cafeteria cost to be allocated- Number of Employees % (170,000.00)           153,000.00         17,000.00                        -  
Total cost after allocation                        -                       -         1,747,000.00       280,000.00 2,027,000.00

> Number 1 is correct! Thank you!…
Number 3, it’s allocation total is correct but the allocated costs are wrong, which leads me to believe that it’s not the method they ask (the weighted average method)…
As for the other questions, I have no idea

Vianney Mon, Oct 18, 2021 12:53 AM